# thanks: Corresponding author

### v1

@@ Abstract
We investigate signatures of quantum chaos and complexity in the quantum annealing Ising model on random Erdős-Rényi graphs. By tuning the connectivity of the graph, the dynamics can be driven from a localized phase through a chaotic regime to an integrable limit. While this dynamical transition reflects in the spectral characteristics, we pursue a broader suite of quantum chaos indicators, some of which can be measured on near-term quantum devices. We study deep thermalization of a quantum state ensemble obtained from a natural unraveling of the subsystem density matrix as an indicator of chaotic dynamics. This extends the analysis of quantum chaos to the ensemble of quantum states. Furthermore, we analyze the eigenstate and eigenvalue correlations through the partial spectral form factor of subsystems and observe distinct signatures of the onset of chaos and its system size dependence, providing experimentally measurable indicators of the localization-to-chaos transition. As a locality-independent probe, we show that the Krylov complexity of operators is also maximized in the chaotic regime, providing a link between graph topology and information scrambling. Finally, we investigate a quantum analogue of the Mpemba effect, where initially “hotter” states can thermalize anomalously fast, a phenomenon most cleanly observed within the chaotic phase. However, away from the chaotic regime, the system is distinguished by multiple crossings across connectivity in its distance from the thermal ensemble with time. Collectively, this work presents a broad characterization of chaos, providing insight beyond spectral analysis and practical indicators for benchmarking near-term quantum devices.
@@ I Introduction
The study of the quantum Ising model has played a central role in modern quantum many-body physics, with implications spanning diverse themes like optimization [1, 2, 3], quantum phase transitions [4], and glassy physics [5, 6, 7]. These models have found practical applications in quantum optimization through quantum annealing algorithms (QAA), which leverage quantum fluctuations to solve challenging optimization problems [8, 9, 10, 11] in noisy intermediate-scale quantum devices. Modern quantum devices exhibit complex graph topologies that emerge from two complementary origins. First, in the encoding of general optimization problems as Ising Hamiltonians [12, 13, 14], the problem structure dictates the qubit interactions. The second stems from hardware constraints in superconducting quantum processors [15, 16], which impose sparse connectivity patterns that resemble semi-random graphs. In such devices, operator spreading [17, 18, 19] and information scrambling [20, 21, 22, 23, 24] correlate with their performance [25, 26, 27]. Operator spreading and information scrambling have been investigated in regular lattice systems with disorder in the context of localization-delocalization transitions [19, 22, 28, 29, 30, 31], but studies on irregular graph topologies are limited. This, therefore, calls for a deeper understanding of how graph topology affects quantum many-body dynamics and the dynamical phases in such devices.
The QAA Hamiltonian generally encodes an Ising-type cost function, where the relative strength of interactions is tuned with time. The connectivity structure of the underlying network is reflected in the spectral properties and dynamics of the quantum Ising system, which serves as the instantaneous QAA Hamiltonian. While sparse connectivity can lead to localized phases, intermediate connectivity regimes exhibit signatures of quantum chaos [32]. In the chaotic regime, the energy level-statistics follow the predictions of random matrix theory (RMT) [33] and the eigenstate thermalization hypothesis (ETH) [34, 35, 36] predicts the eventual approach to the Gibbs ensemble at thermal equilibrium. In the limit of all-to-all connection (in the absence of longitudinal field), the system becomes an integrable Lipkin-Meshkov-Glick (LMG) model [37] exhibiting Poisson level-statistics and breakdown of quantum chaos.
Recent theoretical and experimental developments have extended our understanding of quantum chaos beyond spectral statistics to encompass higher-order correlations in quantum state ensembles, dynamics of operator spreading, and the emergence of complexity under unitary time evolution. In this work, we present an extensive investigation of these aspects in the quantum annealing Ising model on random Erdős-Rényi graphs. We parameterize the connectivity of the graphs with connectance ($\tilde{M}$), which is the connected edges as a fraction of all possible edges. Here, the disorder arises solely from the connectivity pattern instead of interaction strength or external field.
Conventional studies on quantum thermalization focus on the density matrix and the expectation values of local observables at equilibrium predicted by ETH [34, 35, 36]. The recently introduced projected ensemble (PE) [38, 39] provides a natural unraveling of the density matrix to an ensemble of pure quantum states. This allows analysis at the level of the ensemble of states rather than its first moment alone (which is the density matrix). For systems without any non-trivial conserved charges, the PE constructed from the time-evolved state is expected to converge to the Haar ensemble. We obtain the trace distance of the PE from the Haar ensemble as a function of the the system size and $\tilde{M}$ and show that the trace distance shows the fastest decay with the system size at intermediate $\tilde{M}$, indicating the convergence of PE to the Haar ensemble in the chaotic regime. At both the sparse and dense limits ($\tilde{M}\to 0,1$), the trace distance shows a very slow decay with system size. Next, we focus on the partial spectral form factor (pSFF), which goes beyond the spectral form factor (SFF) to capture the correlation among both eigenstates and eigenvalues in the subsystem. PE and pSFF have been demonstrated on multiple platforms [40, 41, 42, 39, 43], motivating the indicators investigated here from an experimental perspective. Thus, they provide a practical benchmarking protocol for quantum annealing devices. To study operator scrambling, we calculate the Krylov complexity (KC) and its late-time saturation with $\tilde{M}$. In the Krylov basis for operators, the initial local operator spreads through a set of Krylov basis states, much like the spread of a particle on a tight-binding model with nearest-neighbor hopping. We observe that the Lanczos coefficients, which are the hopping parameters in the tight-binding notation, display enhanced fluctuations near integrable limits. The KC saturation values are significantly larger in the chaotic regime compared to localized phases. Finally, we investigate the relaxation of simple product states and uncover instances of anomalous thermalization. We observe instances where “hotter” initial states relax more rapidly than “cooler” ones, which is a defining feature of the quantum Mpemba effect (QME). In the chaotic regime, the anomalous relaxation manifests as a smooth, unambiguous decay to equilibrium. This contrasts sharply with the near-integrable and localized limits, where persistent oscillations in the trace distance obscure the relaxation trend. Taken together, this suite of probes provides a detailed picture of the intricate relationship between graph topology and the signatures of quantum thermalization.
@@
This article is organized as follows. Section II provides a brief introduction to the model and the known spectral characteristics. We utilize the projected ensemble protocol to examine deep thermalization through higher-order moments of quantum state ensembles in Sec. III. In Sec. IV, we study the partial spectral form factor (pSFF) of subsystems to identify characteristic signatures of chaotic behavior and its implications in the context of experiments. We also explore the effect of the choice of subsystem on pSFF in this section. Section V presents results on operator delocalization through Krylov complexity. Finally, in Sec. VI we study the relaxation dynamics of initial product states, including the demonstration of a QME. Finally, we conclude in Sec. VII, remarking on the limitations and possible extensions of our study.
II Ising Model on random graph
Figure 1: A single realization of Erdos-Renyi graph with $L=8$ nodes and connectance, $\tilde{M}=0.6$ which implies number of connected edges, $M=17$. Interacting pairs are represented by connected edges.
We consider Ising model as an instantaneous Hamiltonian for the quantum annealing algorithm, which can be expressed as follows,
$H=\frac{J}{M}H_{p}+\frac{g}{L}H_{x}+\frac{h}{L}H_{z}\ ,$ (1)
$H_{x}=-\sum_{i}X_{i},\quad H_{z}=-\sum_{i}Z_{i},\quad H_{p}=-\sum_{i,j}A_{ij}Z_{i}Z_{j}\ ~.$

### v2

@@ Abstract
We investigate signatures of quantum chaos in the mixed-field quantum Ising model on finite-size Erdős-Rényi graphs using probes scalable on near-term quantum devices. By tuning the graph connectivity, the system exhibits a crossover from a localized regime at low connectivity, through a chaotic regime at intermediate connectivity, to a permutation-symmetric integrable limit near all-to-all connectivity. This crossover has possible implications for the performance and trainability of variational algorithms such as QAOA. We characterize this crossover using complementary probes. First, deep thermalization of a projected ensemble starting from a product state reveals slow (fast) convergence to the Haar ensemble at extremal (intermediate) connectivities. Secondly, we analyze eigenstate and eigenvalue correlations using the partial spectral form factor, an experimentally scalable proxy for the spectral form factor with reduced resource overhead, and observe characteristic chaotic signatures at intermediate connectivities and distinct deviations at extremal connectivities. Finally, we explore the Krylov complexity of operators, a locality-independent diagnostic that, although not directly experimentally accessible, serves as a tool for quantifying scrambling. We show that it is maximized deep in the chaotic regime, corroborating the signatures observed through the experimentally scalable probes. Our results provide finite-size benchmarks demonstrating robust signatures of chaos in scalable probes and suggest that these diagnostics can be implemented in current quantum platforms to access regimes beyond classical simulation.
@@ I Introduction
Quantum Ising models have played a central role in the many-body physics of quantum phase transitions [87, 15, 35], arising naturally in both lattice systems and disordered, glassy settings [92, 14, 89]. They also serve as the framework for quantum annealing (QAA) and variational quantum algorithms (VQA), which encode classical optimization problems in an Ising Hamiltonian and use quantum fluctuations to search for low-energy solutions [1, 107, 65, 53, 39, 32, 45]. Hardware connectivity constraints [43, 60] as well as the need to encode general graph optimization problems [79, 47, 46] make the dynamics of these models on random graphs a question of direct practical relevance. Performance of related optimizer circuits [20, 90, 50] has been shown, in specific instances, to correlate with chaos diagnostics such as operator spreading  [18, 94, 17] and information scrambling [91, 68, 48, 51, 98, 76, 72]. While these characteristics have been investigated extensively in regular lattice systems [17, 48, 67, 57, 54, 88], studies on irregular graphs remains largely unexplored.
In QAA, the instantaneous Hamiltonian interpolating between a transverse-field driver and an Ising problem Hamiltonian encoded on a random graph can exhibit chaotic dynamics for initial states with finite excitation energy density. While ideal adiabatic evolution remains confined to the low-energy sector and is therefore insensitive to bulk spectral statistics, it has been shown that adding a chaotic perturbation at intermediate times can improve QAA performance by enhancing the bottleneck gap and delocalizing low-energy states in hard optimization instances [90, 20].
An alternate optimization scheme is the Quantum Approximate Optimization Algorithm (QAOA) [40, 107] which resembles a Trotterization of the QAA. The state is evolved under an alternating sequence of transverse field ($H_{\rm x}$) and cost Hamiltonian ($H_{\rm c}$), with $\{\gamma_{k},\beta_{k}\}$ as variational parameters, optimized to minimize $\langle H_{\rm c}\rangle$ on the final state. Unlike QAA, the intermediate states of a QAOA circuit are not constrained to the low-energy sector. Consequently, eigenstates drawn from the bulk of the spectrum of $H_{\rm c}$ or $H_{\rm x}$ may be transiently populated, making bulk spectral properties directly relevant to the circuit dynamics. In simulations of randomly picked finite-size problem instances, we show (Sec. II) that a three-term ansatz with $H_{\rm c}$, $H_{\rm x}$, and an additional term $H_{\rm d}$ achieves a better approximation ratio than the standard two-term QAOA when $H_{\rm d}$ exhibits signatures of chaos. However, chaotic intermediate dynamics can also be detrimental: they may exacerbate barren plateaus by suppressing parameter gradients at shallow circuit depths [56, 71, 102]. Quantum chaos has also been identified as a key resource in quantum reservoir computing. In an random-graph Ising setup closely related to the one studied here, it was found that optimal learning performance of a quantum reservoir is achieved at the boundary between the chaotic and localized phases [50]. Thus, a refined understanding and control of the onset of chaos in finite-size random-graph Hamiltonians can provide valuable insights into the development of various proposed use cases for NISQ devices.
In this work, we present an investigation of quantum dynamics in the Ising model on Erdős-Rényi (ER) graphs with a mixed field, focusing on how graph connectivity shapes the onset of quantum chaos. We parameterize graph connectivity using the connectance ($\tilde{M}$), which is the fraction of connected edges among all possible edges. For simplicity, we consider the case where Ising coupling strengths are held fixed; disorder in our model arises from the connectivity pattern and not from randomness in the coupling strengths or the external field. Level spacing distributions, level velocities, and spectral correlations in this model were studied in Ref. [42], where chaotic spectral properties were found to emerge in the intermediate $\tilde{M}$ regime. These are spectral diagnostics as they characterize quantum chaos through eigenvalue correlations alone, and carry no direct information about eigenvector structure. et dynamical processes that are experimentally accessible, such as information scrambling, operator growth, and thermalization, are sensitive to eigenvector structure. To build a more complete picture, we employ diagnostics that probe signatures of chaos in projective measurement outcomes, eigenstate structure, operator spreading, and the growth of complexity under unitary evolution — namely, the projected ensemble statistics, partial spectral form factor, and Krylov complexity. These quantities are more directly connected to dynamical and experimentally accessible observables, and together provide a perspective on quantum chaos that is distinct from, and complementary to, the spectral statistics studied previously.
Conventional studies of quantum thermalization focus on quantities governed by the Eigenstate Thermalization Hypothesis (ETH)-specifically, expectation values of local observables  [33, 93, 85]. The projected ensemble (PE) [31, 28] offers a finer characterization by unraveling the density matrix into an ensemble of pure quantum states, enabling analysis beyond the average captured by the density matrix alone. For systems without conserved charges, the PE of a time-evolved state is expected to converge to the Haar ensemble, a condition known as deep thermalization. We show that the distance between PE and Haar distribution decays faster with the system size at intermediate $\tilde{M}$. In contrast, in sparse and dense limits ($\tilde{M}\to 0,1$), the trace distance shows a slow decay with the system size. Next, we focus on the partial spectral form factor (pSFF), which extends the spectral form factor (SFF) to capture correlations between eigenstates and eigenvalues within the subsystem. We empirically demonstrate distinct signatures in the pSFF across different connectivity regimes. Measurements of PE and pSFF have already been experimentally demonstrated in 5255-25 qubit systems in various platforms [52, 34, 105, 28, 103]. While our numerical studies are restricted to $15$ qubits, quantum devices can be used to extend this investigation to larger systems in the near future.
Heisenberg evolution under a chaotic Hamiltonian drives rapid growth of local operators, both in real space and in operator space [80]. To probe the crossover from an operator-spreading perspective, we compute the Krylov complexity (KC) and its late-time saturation as a function of $\tilde{M}$. In the Krylov subspace, an initial local operator spreads over a set of orthogonal basis states generated by Lanczos tridiagonalization of the Liouvillian superoperator, analogous to a particle propagating on a tight-binding chain with nearest-neighbor hopping. KC quantifies this by measuring the average position of the wave packet in the chain. We observe that the Lanczos coefficients, which correspond to the hopping parameters in the tight-binding interpretation, exhibit enhanced fluctuations near the integrable and localized limits (for finite system sizes). Most importantly, the KC saturation values are significantly larger (and grow faster with system size) in the chaotic regime compared to the localized and integrable regimes.
The crossover from chaotic to integrable and localized regimes is associated with the emergence of approximate conservation laws and local integrals of motion [73, 24, 97]. In the model studied here, this takes two distinct forms: at $\tilde{M}\to 0$, disconnected clusters yield trivial local conserved charges, while at $\tilde{M}=1$, a permutation symmetry fragments the Hilbert space (Sec. II). At intermediate $\tilde{M}$, this fragmentation is weakly broken, leaving projections onto Hilbert space fragments as approximate conservation laws that partially obstruct thermalization.
All three diagnostics used in this work, i.e., the projected ensemble, the partial spectral form factor, and the Krylov complexity, probe how these conservation laws suppress quantum information scrambling, but from complementary vantages. PE deviates from the Haar ensemble in the presence of conserved quantities, instead approaching an ensemble constrained by the charges. This was demonstrated in the case of $U(1)$ charges [69, 70] and kinetic constraints [12]. The universal ensemble approached under dynamics with more complex or approximate charges has not been studied extensively, and is expected to deviate from the Haar ensemble. The pSFF, an experimentally accessible proxy for the SFF, captures the resulting effect in eigenvalue and eigenstate correlations [16, 8]. Krylov complexity tracks the restricted growth of local operators along symmetry-protected directions in operator space, with lower saturation values in integrable settings. A central finding of this work is that all three consistently identify the same connectivity-driven crossover, providing a unified picture of how approximate conservation laws govern the onset of chaos across these complementary facets of quantum dynamics.
@@
This article is organized as follows. Section II provides a brief introduction to the model and its symmetries. We also briefly discuss the QAOA protocol and demonstrate the performance of random graph Ising model as a component in circuit ansatz at various graph connectivity. We utilize the projected ensemble protocol to examine deep thermalization through higher-order moments of quantum state ensembles in Sec. III. In Sec. IV, we study the partial spectral form factor (pSFF) of subsystems to identify characteristic signatures of chaotic behavior and its implications in the context of experiments. We also explore the effect of the choice of subsystem on pSFF. Section V presents results on operator delocalization through Krylov complexity. Finally, Sec. VI discusses the limitations and potential extensions of our study.
II Ising Model on random graphs
In this study, we consider the mixed-field quantum Ising model whose Hamiltonian is given by
$H=\frac{J}{M}H_{\rm p}+\frac{g}{L}H_{\rm x}+\frac{h}{L}H_{\rm z}\ ,$ (1)
$H_{\rm p}=-\sum_{i,j}A_{ij}Z_{i}Z_{j},\quad H_{\rm x}=-\sum_{i}X_{i},\quad H_{\rm z}=-\sum_{i}Z_{i}\ ~.$